Given a 32 bit processor, with 2 MB of physical RAM split into 512 frames. What is the size of the single level page table, in entries, if the maximum addressable virtual memory space is 16MB
Answer #1

Size of physical memory = 2 MB = 2 x 220 Bytes = 221 B

Size of virtual memory = 16 MB = 24 x 220 = 224 B

Physical RAM is divided into 512 frames = 29 frames

Size of one frame = (221 ) / (29) = 212 B = 4 KB

Therefore, Size of frame = size of page = 212 B = 4 KB

Total number of pages = size of virtual memory / page size = (224 / 212) = 212 pages.

Each page has a entry. Since there are 212 pages, we have 212 entries. Each entry contains frame number and some other bits. Since we have 29 frames we need 9 bits to store frame number. Neglecting all other bits, size of page table entry = 9 bits.

So, Size of single level page table = No. of entries x size of each entry

= 212 x 9 bits = 4 x 210 x 9 bits = 36K bits = (36/8) K Bytes = 4.5 KB

Therefore, Size of single level Page Table = 4.5 KB